# Dual topological insulator with mirror-symmetry-protected helical edge states

### v1

@@ III The effective $\bm{k}\cdot\bm{p}$ model
Following the approach adopted by Liu et al. [11], we apply the $\bm{k}\cdot\bm{p}$ method together with perturbation theory to derive an effective low-energy Hamiltonian for the quasi-2D Na2CdSn (for details see SM [76]). Our final $4\times 4$ model describes the four energy bands close to $E_{F}$ around the $\Gamma$ point [$\Gamma_{8}$ and $\Gamma_{9}$ bands in Fig. 1-f)].
@@
where $\bm{k}$ is the crystal momentum and $\epsilon(\bm{k})=\frac{\hbar^{2}k^{2}}{2m_{0}}$, with $k=|\bm{k}|$, $\hbar$ the reduced Planck constant, and $m_{0}$ the bare electron mass. The second term in (1),
@@
The perturbative corrections to the $\bm{k}\cdot\bm{p}$ Hamiltonian are calculated by including the remote energy bands via Löwdin perturbation theory or the folding down method. We note that, even though the Löwdin perturbation and folding down techniques are essentially equivalent, the calculations are performed differently in each case. In the Löwdin approach, we can include all the infinite remote bands in the perturbative corrections of the $\bm{k}\cdot\bm{p}$ Hamiltonian up to second order in the crystal momentum. In the folding down, on the other hand, we are only able to account for a finite number of remote bands.
@@ III.1 Löwdin perturbation
The expressions for the first-order terms are obtained by projecting the total Hamiltonian [Eq. (1)] onto a truncated basis consisting of eigenstates of $\mathcal{H}_{0}$ [Eq. (2)]. The matrix elements of the second-order contributions, on the other hand, are calculated using standard perturbation theory formulas [Eq. (S30) in the SM [76]]. In practice, the matrix representation of the second-order contributions is calculated separately and added to the first-order matrix Hamiltonian,
@@
Table 1: First order coupling parameters of the $\bm{k}\cdot\bm{p}$ Hamiltonian  (6). Left and right columns show the parameters and their expressions, respectively. All parameters are real. $\varepsilon_{ij3}$ is the Levi-Civita (permutation) symbol, and repeated indices are summed over, according to the Einstein convention. For more details on the group theory calculations, see Sec. S-V of SM [76].
@@
Projecting each term of Eq. (1) onto the basis $\{|X\uparrow\rangle$, $|Y\uparrow\rangle$, $|X\downarrow\rangle$, $|Y\downarrow\rangle\}$ and employing group theory analysis to find selection rules for the matrix elements (see Sec. S-V of SM [76]) yields
@@
For the second-order contributions we have
@@
with second-order parameters $A_{1}$, $A_{2}$, $D_{1}$, $B$, $C_{1}$, $C_{2}$, $E_{1}$, $F_{1}$ and $F_{2}$ given in Appendix B. Note that here we have taken into account the influence of all the remote bands up to second order in $\bm{k}$. Equations (8)– (14) are in agreement with our derivation using the theory of invariants (see Appendix C).
The eigenvalues of Hamiltonian (6) at $\bm{k}=\bm{0}$ read $\tilde{E}_{1}=\tilde{E}_{3}=C+M$ and $\tilde{E}_{2}=\tilde{E}_{4}=C-M$, with $C=\epsilon_{0}+C_{1}$ and $M=\alpha+C_{2}$. The corresponding eigenvectors are given by $|u_{1(2),\bm{0}}\uparrow\rangle$ and $|u_{3(4),\bm{0}}\downarrow\rangle$ in Table 2. Rewriting the full $\bm{k}$-dependent Hamitlonian in the basis $\{|u_{1,\bm{0}}\uparrow\rangle$, $|u_{2,\bm{0}}\uparrow\rangle$, $|u_{3,\bm{0}}\downarrow\rangle$, $|u_{4,\bm{0}}\downarrow\rangle\}$ we have the final form of our $4\times 4$ effective model
@@
Table 2: Truncated set of zone-center wave functions used as the basis of Hamiltonian (15).
@@
The matrix Hamiltonian in Eq. (15) is block-diagonal, with each block corresponding to a different spin. This is in total agreement with the derivation via folding down (Sec. S-V of SM [76]). Since there is no coupling between these two blocks, the $z$ component of spin is a good quantum number. As we shall see in Subsec. III.3, the parameter $B$, arising from the $H_{\bm{k\cdot p},\bm{k}so}^{(2)}$ and $H_{\bm{k}so,\bm{k\cdot p}}^{(2)}$ terms often neglected for other materials in the literature, is very important for the topology of our system.
Our $\bm{k}\cdot\bm{p}$ Hamiltonian contains off-diagonal matrix elements $Gk_{\pm}^{2}$ (absent in the model of Mao et al.), which are captured by keeping all contributions up to second order in $\bm{k}$ in the Löwdin perturbation theory, folding down technique or theory of invariants (Appendix C). In addition, the Hamiltonian proposed by Mao et al. has terms linear in $k_{\pm}$ in the secondary diagonal, coupling opposite spins. In our model, any term connecting the two spin subspaces is forbidden by symmetry.
@@ III.2 Symmetry analysis
The TRS operator $\mathcal{T}=-i\hat{\sigma}_{y}K$, with $K$ the complex conjugate operator, projected onto the basis $\left\{|u_{1,\bm{0}}\uparrow\rangle,|u_{2,\bm{0}}\uparrow\rangle,|u_{3,\bm{0}}\downarrow\rangle,|u_{4,\bm{0}}\downarrow\rangle\right\}$ (Table 2) is given by $T=-i\sigma_{2}\otimes\tau_{0}K$. We have verified that our model Hamiltonian (15) respects the condition of invariance under TRS by satisfying $TH\left(\bm{k}\right)T^{-1}=H\left(-\bm{k}\right)$.
@@
Hence the eigenvalues of $\hat{\sigma}_{h}$ are $i$ and $-i$. For a 2D crystal invariant under mirror symmetry about the $xy$-plane, $[\mathcal{H},\hat{\sigma}_{h}]=0$, the eigenstates of the Bloch Hamiltonian can be chosen to be also eigenstates of $\hat{\sigma}_{h}$ for all crystal momenta $\bm{k}$. This yields two classes of Bloch eigenstates with mirror eigenvalues $\eta=\pm i$, important for the calculation of the mirror Chern number (Sec. IV).
@@
The model Hamiltonian (15) satisfies [76]
@@
where $u_{\sigma_{h}}$ is the matrix representation of the mirror operator acting on the coordinate space. Therefore the spin texture of the simultaneous eigenstates of $H(\bm{k})$ and $\sigma_{h}$ must be polarized along the $z$ direction in the whole 2D BZ, except at degeneracies [see Fig. 1-g) and Appendix A] [54]. An equation analogous to (21) holds for the remaining symmetry operations of the point group $D_{3h}$ (see SM [76]).
@@ III.3 Fitting
The four energy bands obtained by diagonalizing Hamiltonian (15) are given by
@@
Figure 2: (Color online) Effective energy bands for Na2CdSn. a) Fitting of the $4\times 4$ model Hamiltonian derived via $\bm{k}\cdot\bm{p}$ theory to the DFT energy bands. A range of $\sim 5.5\%$ of the $\Gamma-M$ line was used in the fitting, whose numerical values of the parameters are given in Table 3. Solid (dashed) curves represent the energy bands from the effective model (DFT calculations), and the green (magenta) color represents the $\Gamma-M$ ($\Gamma-K$) high-symmetry line. (b) Spin-up bands with “orbital” (pseudospin) composition. The color bar shows the contribution from the $|X\rangle+i|Y\rangle$ (green) and $|X\rangle-i|Y\rangle$ (brownish red) states. (c) The four energy bands of the model (both spins) with spin expectation value. The spin-up bands are colored red, while the spin-down bands are blue. The bands are spin-degenerate along the $\Gamma-M$ high-symmetry line. In this case, the spin expectation value is not univocally defined. For this reason, we color the bands black in this direction. Along $\Gamma-K$, there is a spin-splitting (not distinguishable in the main graph) due to SOC, $\Delta E_{s}$, which is presented in the inset as a function of the crystal momentum.
In Table 3, we also show the corresponding Hamiltonian terms that contribute to each of the model parameters (see Appendix B for details). For $G=0$, Hamiltonian (15) resembles the well-known Bernevig-Hughes-Zhang (BHZ) model [5]. In both models, the energy bands are inverted if $MB>0$, and therefore, the parameter $B$ is important for the topological analysis of the band structures. We notice that the contributions to the parameter $B$, accounting for all the remote bands perturbatively, come from $H_{\bm{k\cdot p},\bm{k}so}^{(2)}+H_{\bm{k}so,\bm{k\cdot p}}^{(2)}$, i.e., the second-order contribution via $\mathcal{H}_{\bm{k\cdot p}}$ and $\mathcal{H}_{\bm{k}so}$. This suggests that the second-order correction coming from the $\bm{k}$-dependent SOC should not be neglected for Na2CdSn.
@@
In Fig. 2-a), we show the energy bands using the parameters obtained from the fitting (solid curves) and from DFT calculations (dashed). The color code represents the direction in the BZ, with green and magenta for the high-symmetry lines $\Gamma-M$ and $\Gamma-K$, respectively. As expected, our $\bm{k}\cdot\bm{p}$ model reproduces well the DFT data around the $\Gamma$ point but deviates from the numerical data for larger values of the crystal momentum. In addition, note that the energy bands of the analytical model deviate from the DFT bands before they reach the minimum and maximum of the conduction and valence bands, respectively. As a result, the energy gap for the fitted parameters, $E_{g}=285$ meV, is $\sim 21\%$ larger than the gap obtained via DFT calculations. This, however, does not affect the topological properties of the material (see Sec. IV).
In Fig. 2-b), we have the spin-up conduction and valence bands [$E_{1,\bm{k}}$ and $E_{2,\bm{k}}$ in Eqs. (22) and (23), respectively] for the fitted parameters with the “orbital” composition (color code). Note that the conduction band is dominated by the $|X\rangle-i|Y\rangle$ state (brownish red) at the $\Gamma$ point, while the valence band is dominated by $|X\rangle+i|Y\rangle$ (green). For larger values of the crystal momentum, both states become relevant to the two energy bands.
In Fig. 2-c), we show the four energy bands for the fitted parameters, colored according to their spin texture. The bands are spin-degenerate along the $\Gamma-M$ direction, in agreement with the DFT results (see Fig. 1). The spin-splitting along the $\Gamma-K$ line, $\Delta E_{s}\equiv E_{3,\bm{k}}-E_{1,\bm{k}}=E_{4,\bm{k}}-E_{2,\bm{k}}$, is very small, reaching a maximum of 333\,3meV in the range shown. In the inset we display $\Delta E_{s}$ as a function of $\bm{k}$. Note that this spin-splitting is not accompanied by a mixing of the spin components since there is no coupling between spin-up and spin-down blocks in the effective Hamiltonian (15). According to Eq. (24), such a spin-splitting requires that both parameters $A$ and $G$ be finite.
In the next section, we perform a topological analysis of our effective model and predict that the Na2CdSn is a DTI characterized by non-trivial topological invariants.
@@ IV Topology analysis
with $h(\bm{k})=d_{0}(\bm{k})\bar{\tau}_{0}+\bm{d}(\bm{k})\cdot\bm{\bar{\tau}}$. The matrices $\bar{\tau}_{0}$ and $\bm{\bar{\tau}}=\left(\bar{\tau}_{1},\bar{\tau}_{2},\bar{\tau}_{3}\right)$ are the identity and Pauli matrices in the basis $\{i\left(|X\rangle-i|Y\rangle\right)/2,-i\left(|X\rangle+i|Y\rangle\right)/2\}$ of the orbital subspace (pseudospin degree of freedom), respectively. The function $d_{0}(\bm{k})$ and the components of the pseudospin vector $\bm{d}(\bm{k})=\left[d_{1}(\bm{k}),d_{2}(\bm{k}),d_{3}(\bm{k})\right]$ are given by $d_{0}(\bm{k})=C-Dk^{2}$, $d_{1}(\bm{k})=-Ak_{y}+G(k_{x}^{2}-k_{y}^{2})$, $d_{2}(\bm{k})=Ak_{x}-2Gk_{x}k_{y}$ and $d_{3}(\bm{k})=M-Bk^{2}$. The spin-up and -down blocks [$h(\bm{k})$ and $h^{*}(-\bm{k})$, respectively] are connected by TRS.
To investigate in more detail the topology of the Na2CdSn, we perform a square lattice regularization on Eq. (25), which consists of substituting $k_{i}\rightarrow\sin\left(ak_{i}\right)/a$ and $k_{i}^{2}\rightarrow 2\left[1-\cos\left(ak_{i}\right)\right]/a^{2}$ ($i=x,y$) [55]. This substitution is valid for $|k_{i}|\ll 1/a$, and therefore, expected to be valid in our region of interest.
Using the parameters in Table 3, we show in Fig. 3-a) the spin-up energy bands along the high-symmetry path $Y-\Gamma-X$ for both continuous and regularized models. Here, $\Gamma=(0,0)$, $X=(\pi/a,0)$, $Y=(0,\pi/a)$ and $M=(\pi/a,\pi/a)$ are high-symmetry points of the first BZ for the square lattice. The energy bands are plotted only within the region of interest, approximately 3% of the high-symmetry lines, i.e., $0\leq|k|\lesssim 0.09\,a^{-1}$. The black solid and red dashed curves represent the energy bands of the continuous and regularized models, respectively. In Fig. 3-b) we show the spin-up energy bands along $3\%$ of the $M-\Gamma-X$ high-symmetry path, which for the $\Gamma-M$ direction corresponds to $0\leq|k|\lesssim 0.13\,a^{-1}$.
Figure 3: (Color online) Comparison between spin-up energy bands for the continuous (solid black curves) and regularized (dashed red curves) models along $3\%$ of the a) $Y-\Gamma-X$ and b) $M-\Gamma-X$ high-symmetry paths of the square-lattice BZ, corresponding to the region of interest. The crystal momentum intervals along each direction are given by $0\leq|k|<0.09/a$ along either $\Gamma-X$ or $\Gamma-Y$, and $0\leq|k|<0.03\sqrt{2}\pi/a$ along the $\Gamma-M$ direction. c) Pseudospin skyrmion configuration of the regularized model over the first BZ, i.e., $0<|k_{x(y)}|<\pi/a$. Each arrow is determined by the components of the pseudospin unit vector at the corresponding $\bm{k}$ point, $\hat{\bm{d}}(\bm{k})$, with the color code given by the $\hat{d}_{3}(\bm{k})$ component. The inset shows a zoom into the region of interest, where we can see in more detail the behavior of the vector field around the $\Gamma$ point. The ranges for the components of the crystal momentum in the inset are given by $0<|k_{x(y)}|<0.02\pi/a$.
@@
The pseudospin unit vector for the regularized model, $\hat{\bm{d}}(\bm{k})=\bm{d}(\bm{k})/|\bm{d}(\bm{k})|$, defines a mapping from the compact BZ to the unit Bloch sphere. The number of times this mapping wraps the Bloch sphere defines an integer-quantized topological invariant known as the winding number [83, 84, 55]. In Fig. 3-c), we show the skyrmion configuration obtained for Na2CdSn. Each vector is defined by the three components of $\hat{\bm{d}}(\bm{k})$ at the corresponding $\bm{k}$ point of the BZ, and the color code denotes the $\hat{d}_{3}(\bm{k})$ component. We can see that at the $\Gamma$ point the vector field points in the positive $z$ direction, while at the corners of the BZ, the vectors point in the negative $z$ direction. Additionally, the skyrmion configuration in Fig. 3-c) resembles that of a second-order antiskyrmion, which alludes to a topological charge of $-2$ [85, 86, 87, 88].
To unequivocally identify the topological phase of the regularized model, we calculate the Chern number for the valence bands of each spin sector separately. The first step is to define the Berry connection, $\bm{A}_{v}^{\uparrow(\downarrow)}(\bm{k})=-i\langle u_{v}^{\uparrow(\downarrow)}(\bm{k})|\bm{\nabla}_{\bm{k}}|u_{v}^{\uparrow(\downarrow)}(\bm{k})\rangle$, and Berry curvature, $F^{\uparrow(\downarrow)}_{v}(\bm{k})=\hat{z}\cdot\left[\bm{\nabla}_{\bm{k}}\times\bm{A}_{v}^{\uparrow(\downarrow)}(\bm{k})\right]$ [55], where $|u_{v}^{\uparrow(\downarrow)}(\bm{k})\rangle$ denote the valence band of the corresponding spin sector. Now, we define the Chern number for each spin sector using the well-known TKNN formula [89], which consists of integrating the Berry curvature over the whole BZ, i.e. [55]

### v2

@@ III The effective $\bm{k}\cdot\bm{p}$ model
Following the approach adopted by Liu et al. [11], we apply the $\bm{k}\cdot\bm{p}$ method together with perturbation theory to derive an effective low-energy Hamiltonian for the quasi-2D Na2CdSn (for details, see SM [75]). Our final $4\times 4$ model describes the four energy bands close to $E_{F}$ around the $\Gamma$ point [$\Gamma_{8}$ and $\Gamma_{9}$ bands in Fig. 1-f)].
@@
where $\bm{k}$ is the crystal momentum and $\epsilon(\bm{k})=\frac{\hbar^{2}k^{2}}{2m_{0}}$, with $k=|\bm{k}|$, $\hbar$ the reduced Planck constant, and $m_{0}$ the bare electron mass. The second term in Eq. (1),
@@
The perturbative corrections to the $\bm{k}\cdot\bm{p}$ Hamiltonian are calculated by including the remote energy bands via Löwdin perturbation theory or the folding-down method. We note that, even though the Löwdin perturbation and folding-down techniques are essentially equivalent, the calculations are performed differently in each case. In the Löwdin approach, we can include all the infinite remote bands in the perturbative corrections of the $\bm{k}\cdot\bm{p}$ Hamiltonian up to second order in the crystal momentum. In the folding down, on the other hand, we can only account for a finite number of remote bands.
@@ III.1 Löwdin perturbation
The expressions for the first-order terms are obtained by projecting the total Hamiltonian [Eq. (1)] onto a truncated basis consisting of eigenstates of $\mathcal{H}_{0}$ [Eq. (2)]. The matrix elements of the second-order contributions, on the other hand, are calculated using standard perturbation theory formulas (Eq. (S30) in the SM [75]). In practice, the matrix representation of the second-order contributions is calculated separately and added to the first-order matrix Hamiltonian:
@@
Table 1: First-order coupling parameters of the $\bm{k}\cdot\bm{p}$ Hamiltonian in Eq.  (6). Left and right columns show the parameters and their expressions, respectively. All parameters are real. $\varepsilon_{ij3}$ is the Levi-Civita (permutation) symbol, and repeated indices are summed over, according to the Einstein convention. For more details on the group theory calculations, see Sec. S-V of SM [75].
@@
Projecting each term of Eq. (1) onto the basis $\{|X\uparrow\rangle$, $|Y\uparrow\rangle$, $|X\downarrow\rangle$, $|Y\downarrow\rangle\}$ and employing group theory analysis to find selection rules for the matrix elements (see Sec. S-V in the SM [75]) yields
@@
For the second-order contributions, we have
@@
with second-order parameters $A_{1}$, $A_{2}$, $D_{1}$, $B$, $C_{1}$, $C_{2}$, $E_{1}$, $F_{1}$ and $F_{2}$ given in Appendix B. Note that, here, we have considered the influence of all the remote bands up to second order in $\bm{k}$. Equations (8)–(14) are in agreement with our derivation using the theory of invariants (see Appendix C).
The eigenvalues of the Hamiltonian in Eq. (6) at $\bm{k}=\bm{0}$ read $\tilde{E}_{1}=\tilde{E}_{3}=C+M$ and $\tilde{E}_{2}=\tilde{E}_{4}=C-M$, with $C=\epsilon_{0}+C_{1}$ and $M=\alpha+C_{2}$. The corresponding eigenvectors are given by $|u_{1(2),\bm{0}}\uparrow\rangle$ and $|u_{3(4),\bm{0}}\downarrow\rangle$ in Table 2. Next, we follow a usual procedure in the $\bm{k}\cdot\bm{p}$ literature [49, 51, 52] and perform a change of basis from $\{|X\uparrow\rangle$, $|Y\uparrow\rangle$, $|X\downarrow\rangle$, $|Y\downarrow\rangle\}$ to $\{|u_{1,\bm{0}}\uparrow\rangle$, $|u_{2,\bm{0}}\uparrow\rangle$, $|u_{3,\bm{0}}\downarrow\rangle$, $|u_{4,\bm{0}}\downarrow\rangle\}$, rewriting the full $\bm{k}$-dependent Hamiltonian in a simpler form (diagonal at $\bm{k}=0$):
@@
Table 2: Truncated set of zone-center wave functions used as the basis of the Hamiltonian in Eq. (15).
@@
The matrix Hamiltonian in Eq. (15) is block diagonal, with each block corresponding to a different spin. This is in total agreement with the derivation via folding down (Sec. S-V in the SM [75]). Since there is no coupling between these two blocks, the $z$ component of spin is a good quantum number. As we shall see in Sec. III.3, the parameter $B$, arising from the $H_{\bm{k\cdot p},\bm{k}so}^{(2)}$ and $H_{\bm{k}so,\bm{k\cdot p}}^{(2)}$ terms often neglected for other materials in the literature, is very important for the topology of our system.
Our $\bm{k}\cdot\bm{p}$ Hamiltonian contains off-diagonal matrix elements $Gk_{\pm}^{2}$ (absent in the model of Mao et al. [70]), which are captured by keeping all contributions up to second order in $\bm{k}$ in the Löwdin perturbation theory, folding-down technique, or theory of invariants (Appendix C). In addition, the Hamiltonian proposed by Mao et al. [70] has terms linear in $k_{\pm}$ in the secondary diagonal, coupling opposite spins. In our model, any term connecting the two spin subspaces is forbidden by symmetry  [^2] By applying the theory of invariants using an extended basis set, we have verified that the next remote band, $\Gamma_{4}$, couples to the bands of interest, $\Gamma_{6}$, through zeroth- and first-order in $\bm{k}$ spin-mixing terms. However, after eliminating the remote bands via Löwdin perturbation theory, the final $4\times 4$ Hamiltonian remains equivalent to Eq. (15), with the new terms merely renormalising the model parameters. This result is in line with our model derivation in Appendix C, where we have obtained all terms allowed by symmetry through the theory of invariants.. This result was obtained via all the three perturbative approaches considered in this paper.
@@ III.2 Symmetry analysis
The TRS operator $\mathcal{T}=-i\hat{\sigma}_{y}K$, with $K$ the complex conjugate operator, projected onto the basis $\left\{|u_{1,\bm{0}}\uparrow\rangle,|u_{2,\bm{0}}\uparrow\rangle,|u_{3,\bm{0}}\downarrow\rangle,|u_{4,\bm{0}}\downarrow\rangle\right\}$ (Table 2) is given by $T=-i\sigma_{2}\otimes\tau_{0}K$. We have verified that our model Hamiltonian in Eq. (15) respects the condition of invariance under TRS by satisfying $TH\left(\bm{k}\right)T^{-1}=H\left(-\bm{k}\right)$.
@@
Hence the eigenvalues of $\hat{\sigma}_{h}$ are $i$ and $-i$. For a 2D crystal invariant under mirror symmetry about the $xy$-plane, $[\mathcal{H},\hat{\sigma}_{h}]=0$, the eigenstates of the Bloch Hamiltonian can be chosen to also be eigenstates of $\hat{\sigma}_{h}$ for all crystal momenta $\bm{k}$. This yields two classes of Bloch eigenstates with mirror eigenvalues $\eta=\pm i$, important for the calculation of the mirror Chern number (Sec. IV).
@@
The model Hamiltonian in Eq. (15) satisfies [75]
@@
where $u_{\sigma_{h}}$ is the matrix representation of the mirror operator acting on the coordinate space. Therefore the spin texture of the simultaneous eigenstates of $H(\bm{k})$ and $\sigma_{h}$ must be polarized along the $z$ direction in the whole 2D BZ, except at degeneracies [see Fig. 1-g) and Appendix A] [54]. An equation analogous to Eq. (21) holds for the remaining symmetry operations of the point group $D_{3h}$ (see SM [75]).
@@ III.3 Fitting
The four energy bands obtained by diagonalizing the Hamiltonian in Eq. (15) are given by
@@
Figure 2: (Color online) Effective energy bands for Na2CdSn. a) Fitting of the $4\times 4$ model Hamiltonian derived via $\bm{k}\cdot\bm{p}$ theory to the density functional theory (DFT) energy bands. A range of $\sim 5.5\%$ of the $\Gamma-M$ line was used in the fitting, whose numerical values of the parameters are given in Table 3. Solid (dashed) curves represent the energy bands from the effective model (DFT calculations), and green (magenta) represents the $\Gamma-M$ ($\Gamma-K$) high-symmetry line. (b) Spin-up bands with orbital (pseudospin) composition. The color bar shows the contribution from the $|X\rangle+i|Y\rangle$ (green) and $|X\rangle-i|Y\rangle$ (brownish red) states. (c) The four energy bands of the model (both spins) with spin expectation value. The spin-up bands are colored red, while the spin-down bands are blue. The bands are spin degenerate along the $\Gamma-M$ high-symmetry line. In this case, the spin expectation value is not univocally defined. For this reason, we color the bands black in this direction. Along $\Gamma-K$, there is a spin splitting (not distinguishable in the main graph) due to spin-orbit coupling (SOC), $\Delta E_{s}$, which is presented in the inset as a function of the crystal momentum.
In Table 3, we also show the corresponding Hamiltonian terms that contribute to each of the model parameters (see Appendix B for details). For $G=0$, the Hamiltonian in Eq. (15) resembles the well-known Bernevig-Hughes-Zhang (BHZ) model [5]. In both models, the energy bands are inverted if $MB>0$, and therefore, the parameter $B$ is important for the topological analysis of the band structures. We notice that the contributions to the parameter $B$, accounting for all the remote bands perturbatively, come from $H_{\bm{k\cdot p},\bm{k}so}^{(2)}+H_{\bm{k}so,\bm{k\cdot p}}^{(2)}$, i.e., the second-order contribution via $\mathcal{H}_{\bm{k\cdot p}}$ and $\mathcal{H}_{\bm{k}so}$. This suggests that the second-order correction coming from the $\bm{k}$-dependent SOC should not be neglected for Na2CdSn.
@@
In Fig. 2-a), we show the energy bands using the parameters obtained from the fitting (solid curves) and from DFT calculations (dashed). The color code represents the direction in the BZ, with green and magenta for the high-symmetry lines $\Gamma-M$ and $\Gamma-K$, respectively. As expected, our $\bm{k}\cdot\bm{p}$ model reproduces well the DFT data around the $\Gamma$ point but deviates from the numerical data for larger values of the crystal momentum. In addition, note that the energy bands of the analytical model deviate from the DFT bands before they reach the minimum and maximum of the conduction and valence bands, respectively. As a result, the energy gap for the fitted parameters $E_{g}=285$ meV is $\sim 21\%$ larger than the gap obtained via DFT calculations. This, however, does not affect the topological properties of the material (see Sec. IV).
In Fig. 2-b), we show the spin-up conduction and valence bands [$E_{1,\bm{k}}$ and $E_{2,\bm{k}}$ in Eqs. (22) and (23), respectively] for the fitted parameters with the orbital composition (color code). Note that the conduction band is dominated by the $|X\rangle-i|Y\rangle$ state (brownish red) at the $\Gamma$ point, while the valence band is dominated by $|X\rangle+i|Y\rangle$ (green). For larger values of the crystal momentum, both states become relevant to the two energy bands.
In Fig. 2-c), we show the four energy bands for the fitted parameters, colored according to their spin texture. The bands are spin degenerate along the $\Gamma-M$ direction, in agreement with the DFT results (see Fig. 1). The spin splitting along the $\Gamma-K$ line, $\Delta E_{s}\equiv E_{3,\bm{k}}-E_{1,\bm{k}}=E_{4,\bm{k}}-E_{2,\bm{k}}$, is very small, reaching a maximum of 333\,3meV in the range shown. In the inset we display $\Delta E_{s}$ as a function of $\bm{k}$. Note that this spin splitting is not accompanied by a mixing of the spin components since there is no coupling between spin-up and spin-down blocks in the effective Hamiltonian in Eq. (15). According to Eq. (24), such a spin splitting requires that both parameters $A$ and $G$ be finite.
In the next section, we perform a topological analysis of our effective model and predict that the Na2CdSn is a DTI characterized by nontrivial topological invariants.
@@ IV Topology analysis
with $h(\bm{k})=d_{0}(\bm{k})\bar{\tau}_{0}+\bm{d}(\bm{k})\cdot\bm{\bar{\tau}}$. The matrices $\bar{\tau}_{0}$ and $\bm{\bar{\tau}}=\left(\bar{\tau}_{1},\bar{\tau}_{2},\bar{\tau}_{3}\right)$ are the identity and Pauli matrices in the basis $\{i\left(|X\rangle-i|Y\rangle\right)/2,-i\left(|X\rangle+i|Y\rangle\right)/2\}$ of the orbital subspace (pseudospin degree of freedom), respectively. The function $d_{0}(\bm{k})$ and the components of the pseudospin vector $\bm{d}(\bm{k})=\left[d_{1}(\bm{k}),d_{2}(\bm{k}),d_{3}(\bm{k})\right]$ are given by $d_{0}(\bm{k})=C-Dk^{2}$, $d_{1}(\bm{k})=-Ak_{y}+G(k_{x}^{2}-k_{y}^{2})$, $d_{2}(\bm{k})=Ak_{x}-2Gk_{x}k_{y}$, and $d_{3}(\bm{k})=M-Bk^{2}$. The spin-up and down blocks [$h(\bm{k})$ and $h^{*}(-\bm{k})$, respectively] are connected by TRS.
To investigate in more detail the topology of Na2CdSn, we perform a square lattice regularization on Eq. (25), which consists of substituting $k_{i}\rightarrow\sin\left(ak_{i}\right)/a$ and $k_{i}^{2}\rightarrow 2\left[1-\cos\left(ak_{i}\right)\right]/a^{2}$ ($i=x,y$) [55]. This substitution is valid for $|k_{i}|\ll 1/a$ and, therefore, expected to be valid in our region of interest.
Using the parameters in Table 3, we show in Fig. 3-a) the spin-up energy bands along the high-symmetry path $Y-\Gamma-X$ for both continuous and regularized models. Here, $\Gamma=(0,0)$, $X=(\pi/a,0)$, $Y=(0,\pi/a)$ and $M=(\pi/a,\pi/a)$ are high-symmetry points of the first BZ for the square lattice. The energy bands are plotted only within the region of interest, $\sim$3of the high-symmetry lines, i.e., $0\leq|k|\lesssim 0.09\,a^{-1}$. The black solid and red dashed curves represent the energy bands of the continuous and regularized models, respectively. In Fig. 3-b) we show the spin-up energy bands along $3\%$ of the $M-\Gamma-X$ high-symmetry path, which for the $\Gamma-M$ direction corresponds to $0\leq|k|\lesssim 0.13\,a^{-1}$.
Figure 3: (Color online) Comparison between spin-up energy bands for the continuous (solid black curves) and regularized (dashed red curves) models along $3\%$ of the a) $Y-\Gamma-X$ and b) $M-\Gamma-X$ high-symmetry paths of the square-lattice Brillouin zone (BZ), corresponding to the region of interest. The crystal momentum intervals along each direction are given by $0\leq|k|<0.09/a$ along either $\Gamma-X$ or $\Gamma-Y$, and $0\leq|k|<0.03\sqrt{2}\pi/a$ along the $\Gamma-M$ direction. c) Pseudospin skyrmion configuration of the regularized model over the first BZ, i.e., $0<|k_{x(y)}|<\pi/a$. Each arrow is determined by the components of the pseudospin unit vector at the corresponding $\bm{k}$ point, $\hat{\bm{d}}(\bm{k})$, with the color code given by the $\hat{d}_{3}(\bm{k})$ component. The inset shows a zoom into the region of interest, where we can see in more detail the behavior of the vector field around the $\Gamma$ point. The ranges for the components of the crystal momentum in the inset are given by $0<|k_{x(y)}|<0.02\pi/a$.
@@
The pseudospin unit vector for the regularized model $\hat{\bm{d}}(\bm{k})=\bm{d}(\bm{k})/|\bm{d}(\bm{k})|$ defines a mapping from the compact BZ to the unit Bloch sphere. The number of times this mapping wraps the Bloch sphere defines an integer-quantized topological invariant known as the winding number [85, 86, 55]. In Fig. 3-c), we show the skyrmion configuration obtained for Na2CdSn. Each vector is defined by the three components of $\hat{\bm{d}}(\bm{k})$ at the corresponding $\bm{k}$ point of the BZ, and the color code denotes the $\hat{d}_{3}(\bm{k})$ component. We can see that, at the $\Gamma$ point, the vector field points in the positive $z$ direction, while at the corners of the BZ, the vectors point in the negative $z$ direction. Additionally, the skyrmion configuration in Fig. 3-c) resembles that of a second-order antiskyrmion, which alludes to a topological charge of $-2$ [87, 88, 89, 90].
To unequivocally identify the topological phase of the regularized model, we calculate the Chern number for the valence bands of each spin sector separately. The first step is to define the Berry connection $\bm{A}_{v}^{\uparrow(\downarrow)}(\bm{k})=-i\langle u_{v}^{\uparrow(\downarrow)}(\bm{k})|\bm{\nabla}_{\bm{k}}|u_{v}^{\uparrow(\downarrow)}(\bm{k})\rangle$ and Berry curvature $F^{\uparrow(\downarrow)}_{v}(\bm{k})=\hat{z}\cdot\left[\bm{\nabla}_{\bm{k}}\times\bm{A}_{v}^{\uparrow(\downarrow)}(\bm{k})\right]$ [55], where $|u_{v}^{\uparrow(\downarrow)}(\bm{k})\rangle$ denotes the valence band of the corresponding spin sector. Now we define the Chern number for each spin sector using the well-known TKNN formula [91], which consists of integrating the Berry curvature over the whole BZ [55]: